# The Rees algebra and analytic spread of a divisorial filtration

### v1

@@ 2. Valuations, Filtrations and Rees Algebras
We will denote the maximal ideal of a local ring $R$ by $m_{R}$. The length of an $R$-module $M$ will be denoted by $\lambda_{R}(M)$ or $\lambda(M)$.
@@ 2.1. Extensions of valuations and Rees algebras
where $m_{v}$ is the maximal ideal of $\mathcal{O}_{v}$. We have that $\sqrt{I(v)_{\lambda}\cap R}=\mbox{center}_{R}(v)$ for all $\lambda\in vK_{>0}$.
@@
Let $K\rightarrow L$ be a finite field extension, $R$ be a normal noetherian domain with quotient field $K$ and $S$ be the integral closure of $R$ in $L$. Suppose that $S$ is a finite $R$-module. Let $v$ be a valuation of $K$ which is nonnegative on $R$, and $w$ be an extension of $v$ to $L$. Then $w$ is nonnegative on $S$ and
@@
The maximal ideal $m_{w}$ of $\mathcal{O}_{w}$ contracts to the maximal ideal $m_{v}$ of $\mathcal{O}_{v}$ so $\mbox{center}_{S}(w)$ contracts to $\mbox{center}_{R}(v)$. Thus the lemma holds by Corollary A.8 [3]. ∎
Let $R$ be a normal semi local domain with quotient field $K$, $v_{1},\ldots,v_{r}$ be rank 1 valuations of $K$ which are nonnegative on $R$ and $\lambda_{i}\in v_{i}K$ for $1\leq i\leq r$. Let $J(R)$ be the intersection of the maximal ideals of $R$. Then
$\left[\cap_{i\in\Lambda}I(v_{i})_{\lambda_{i}}\right]\cap R=\left[\cap_{i=1}^{r}I(v_{i})_{\lambda_{i}}\right]\cap R:J(R)^{\infty}.$
@@
The following lemma follows from Lemma 2.2.
@@
Let $K$ be a field and $v_{1},\ldots,v_{r}$ be rank 1 valuations of $K$ and $\lambda_{i}\geq 0$ be elements of $v_{i}K$ for $1\leq i\leq r$. Let $R$ be a normal semilocal domain with quotient field $K$ such that the $v_{i}$ are nonnegative on $R$ for all $i$. Define a graded filtration $\mathcal{I}(R)=\{I(R)_{n}\}$ of $R$ by $I(R)_{n}=(\cap_{i}I(v_{i})_{n\lambda_{i}})\cap R$. Suppose that $K\rightarrow L$ is a finite field extension. Let $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$ for $1\leq i\leq r$. Let $S$ be the integral closure of $R$ in $L$.
(10) Define a graded filtration $\mathcal{I}(S)=\{I(S)_{n}\}$ on $S$ by $I(S)_{n}=(\cap_{i}\cap_{j}I(w_{ij})_{n\lambda_{i}})\cap S$.
@@
Let $N$ be a finite extension of $L$ such that $N$ is a normal extension of $K$. Let $G=\mbox{Aut}_{K}(N)$. We have a tower of fields $K\rightarrow N^{G}\rightarrow N$ where $N^{G}$ is the fixed field of $N$ under the action of $G$. The extension $K\rightarrow N^{G}$ is purely inseparable and $N^{G}\rightarrow N$ is Galois with Galois group $G$ (c.f. Proposition V.6.11 [25]).
@@
(13) $w_{i,j}=w_{i}\sigma_{\tau_{i}(j)}$
for all $j$ (Corollary VI.7.3 [37]). For $b\in\{1,\ldots,a\}$, we compute
@@
where the minimum is over $1\leq\alpha_{1}\leq\cdots\leq\alpha_{a}\leq b$. Thus
@@ 2.2. Divisorial Filtrations
A divisorial valuation of a domain $R$ ([36, Definition 9.3.1]) is a valuation $\nu$ of $K$ such that if $\mathcal{O}_{\nu}$ is the valuation ring of $\nu$ with maximal ideal $\mathfrak{m}_{\nu}$, then $R\subset\mathcal{O}_{\nu}$ and if $\mathfrak{p}=\mathfrak{m}_{\nu}\cap R$ then $\mbox{trdeg}_{\varkappa(\mathfrak{p})}\varkappa(\nu)={\rm ht}(\mathfrak{p})-1$, where $\varkappa(\mathfrak{p})$ is the residue field of $R_{\mathfrak{p}}$ and $\varkappa(\nu)$ is the residue field of $\mathcal{O}_{\nu}$. If $\nu$ is a divisorial valuation of R such that $m_{R}=m_{\nu}\cap R$, then $\nu$ is called an $m_{R}$-valuation.
@@
By [36, Theorem 9.3.2], the valuation ring of every divisorial valuation $\nu$ is Noetherian, hence is a discrete valuation. Suppose that $R$ is an excellent local domain. Then a valuation $\nu$ of the quotient field $K$ of $R$ which is nonnegative on $R$ is a divisorial valuation of $R$ if and only if the valuation ring $\mathcal{O}_{\nu}$ is essentially of finite type over $R$ ([12, Lemma 6.1]).
A divisorial filtration of an excellent domain $R$ is a filtration =$\mathcal{I}=\{I_{n}\}$ where there exist divisorial valuations $v_{1},\ldots,v_{r}$ of $R$ and $\lambda_{1},\ldots,\lambda_{n}\in{\mathbb{R}}_{\geq 0}$ such that $I_{n}=I(v_{1})_{n\lambda_{1}}\cap\cdots\cap I(v_{r})_{n\lambda_{r}}$. A representation of a divisorial valuation $\mathcal{I}$ is a filtration $\mathcal{I}=\mathcal{I}(D)$ where $D$ is an effective ${\mathbb{R}}$-Weil divisor on a normal $R$-scheme $X$ with a projective birational morphism $\pi:X\rightarrow\mbox{Spec}(R)$ such that $I_{n}=I(nD)=\Gamma(X,\mathcal{O}_{X}(-\lceil nD\rceil))$ for $n\geq 0$. If $D=\sum\lambda_{i}E_{i}$ where $E_{i}$ are prime divisors on $X$, then $\lceil nD\rceil=\sum\lceil n\lambda_{i}\rceil E_{i}$ where $\lceil x\rceil$ is the round up of a real number $x$. $\mathcal{I}(D)$ is called a ${\mathbb{Q}}$-divisorial filtration if $D$ is an effective ${\mathbb{Q}}$-divisor and is called a ${\mathbb{Z}}$-divisorial filtration, or just a divisorial filtration if $D$ is a ${\mathbb{Z}}$-divisor.
@@
Associated to a discrete valuation ring $V$ is a unique canonical valuation $v_{V}$ defined by $v_{V}(x)=n$ if $x\in m_{V}^{n}\setminus m_{V}^{n+1}$ for $x\in V$ and where $m_{V}$ is the maximal ideal of $V$. If $w$ is any valuation whose valuation ring is $V$, so that $w$ is an equivalent valuation to $v$, we have that $w=cv$ where $c=w(t)$ for $t$ a generator of $m_{V}$.
Suppose that =$\mathcal{I}=\{I_{n}\}$ where $I_{n}=I(v_{1})_{n\lambda_{1}}\cap\cdots\cap I(v_{r})_{n\lambda_{r}}$ is a divisorial filtration of $R$ and $D=\sum_{i=1}^{r}\lambda_{i}E_{i}$ is an ${\mathbb{R}}$-Weil divisor on a normal $R$-scheme $X$ and $E_{i}$ are prime divisors on $X$ such that $X$ is projective and birational over $\mbox{Spec}(R)$. Let $v_{1},\ldots,v_{r}$ be the canonical valuations of the valuation rings $\mathcal{O}_{X,E_{i}}$. Then
$\Gamma(X,\mathcal{O}_{X}(-\lceil nD\rceil))=I(v_{1})_{n\lambda_{1}}\cap\cdots\cap I(v_{r})_{n\lambda_{r}}.$
@@
Let $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$. Let $u_{ij}$ be the canonical valuation of $\mathcal{O}_{w_{ij}}$. There exists a prime divisor $F_{ij}$ on $Y$ such that $\mathcal{O}_{w_{ij}}=\mathcal{O}_{F_{ij}}$. Write $w_{ij}=d_{ij}u_{ij}$. Let $D^{\prime}=\sum_{i,j}\frac{\lambda_{i}}{d_{ij}}F_{ij}$. Then
(15) $I(D^{\prime})_{n}=\cap_{i}\cap_{j}I(u_{ij})_{n\frac{\lambda_{i}}{d_{ij}}}=\cap_{i}\cap_{j}I(w_{ij})_{n\lambda_{i}}$
@@
We now weaken our assumptions on $R$ and suppose that $R$ is an arbitrary excellent local ring. Let $P_{1},\ldots,P_{r}$ be the minimal prime ideals of $R$. Let $\varphi_{i}:R\rightarrow R/P_{i}$ be the natural projection for $1\leq i\leq r$. Let $\{v_{ij}\}$ be divisorial valuations on $R/P_{i}$ and $\lambda_{ij}\in{\mathbb{R}}_{\geq 0}$. Let =$\mathcal{I}=\{I_{n}\}$ where
(16) $I_{n}=\cap_{i=1}^{r}\cap_{i,j}\varphi_{i}^{-1}(I(v_{ij})_{n\lambda_{ij}}).$
If $P$ is a prime ideal of a local ring $R$, $\varphi:R\rightarrow R/P$ and $v$ is a valuation of the quotient field of $R/P$, then $w=v\circ\varphi$ is a valuation of $R$, as defined in VI.3.1 [2]. That is, by extending the value group $vK/P$ of $V$ by introducing the element $\infty$, which is larger than any element of $vK/P$, and letting $\Gamma_{\infty}=vK/P\cap\{\infty\}$, the mapping $w:R\rightarrow n\Gamma_{\infty}$ satisfies the properties
(17) $\begin{array}[]{l}v(xy)=v(x)+v(y)\mbox{ for }x,y\in R\\ v(x+y)\geq\min\{v(x),v(y)\}\mbox{ for }x,y\in R\\ v(1)=0\mbox{ and }v(0)=\infty\end{array}$
With the notation of (16), the mappings $w_{ij}\circ\varphi_{i}$ are valuations of $R$, satisfying the conclusions of (17).
Let $R$ be a local ring and $w$ be a valuation of $R$. For $\lambda\geq 0$ in $\Gamma_{\infty}$, let
$I(w)_{\lambda}=\{x\in R\mid v(c)\geq\lambda\},$
an ideal in $R$. The center of $w$ on $R$ is the prime ideal $\mbox{center}_{R}(v)=\{x\in R\mid w(x)>0\}$.
@@ 3. A generalized McAdam theorem
In this section we prove Theorem 1.4, which is stated in the introduction.
@@ 3.1. Resolutions and Alterations
The proof of Section 5 of [11] shows the following.
@@
(Section 5 [11]) Let $R$ be a normal excellent local domain and let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. Suppose that there exists a projective birational morphism $\pi:X\rightarrow\mbox{Spec}(R)$ such that $X$ is nonsingular, and there exists an effective ${\mathbb{Q}}$-divisor $D$ on $X$ such that $\mathcal{I}=\mathcal{I}(D)$. If $m_{R}\in\mbox{Ass}(R/I_{m_{0}})$ for some $m_{0}\in{\mathbb{Z}}_{>0}$, then the analytic spread of $\mathcal{I}$ is $\ell(\mathcal{I})=\dim R$.
The following proposition is an extension of Section 6 of [11], with essentially the same proof.
@@
(Section 6 [11] Let $R\rightarrow S$ be a finite extension of reduced rings where $R$ is a Noetherian local ring of dimension $d$ and $S$ is a semi local ring. Let $\mathcal{J}=\{J_{n}\}$ be a graded filtration of $S$ and let =$\mathcal{I}=\{I_{n}\}$ be the graded filtration of $R$ defined by $I_{n}=J_{n}\cap R$. For $q$ a prime ideal of $S$, let $\mathcal{J}_{q}$ be the graded filtration $\mathcal{J}_{q}=\{(J_{n})_{q}\}$ of $S_{q}$. Suppose that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$ for some $m_{0}>0$. Then there exists a maximal ideal $q$ of $S$ such that $q_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$. If we further have that $R$ and $S$ are equidimensional of dimension $d$, $\ell(\mathcal{J}_{q})=d$ and there exists a nonzero divisor of $R$ in $M=\mbox{Ann}_{R}(S/R)$, then $\ell(\mathcal{I})=d$.
@@
Let $\{\mathfrak{m}_{i}$} be the maximal ideals of $S$. Since we assume that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$, there exists $y\in R/I_{n_{0}}$ such that $m_{R}=\mbox{ann}_{R}(y)$. We have that $y\in S/J_{n}$ is nonzero since $R/I_{m_{0}}\rightarrow S/J_{m_{0}}$ is an inclusion. Thus $\mbox{Ann}_{S}(y)\neq S$. Since maximal elements in the set of annihilators of elements of $S/J_{m_{0}}$ are prime ideals (by [28, Theorem 6.1]), there exists a prime ideal $q$ in $S$ which contains $\mbox{Ann}_{S}(y)$ and is the annihilator of an element $z$ of $S/J_{m_{0}}$. We have that $q$ contains $m_{R}S$ and
@@
so $q$ is a maximal ideal of $S$ such that $q\in\mbox{Ass}_{S}(S/J_{m_{0}})$, and thus $q_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$.
@@
This follows from Remark 6.6 to Lemma 6.5 of [12]. ∎
@@
(Definition 2.20 [22]) Let $X$ be a Noetherian integral scheme. An Alteration $X^{\prime}$ of $X$ is an integral scheme $X^{\prime}$ with a morphism $\varphi:X^{\prime}\rightarrow X$ which is dominant, proper and such that for some open $U\subset X$, $\varphi^{-1}(U)\rightarrow U$ is finite.
A scheme $X$ is nonsingular if $\mathcal{O}_{X,q}$ is a regular local ring for all $q\in X$.

### v2

@@ 2. Valuations, Filtrations and Rees Algebras
A local ring will be assumed to be Noetherian. We will denote the maximal ideal of a local ring $R$ by $m_{R}$. The length of an $R$-module $M$ will be denoted by $\lambda_{R}(M)$ or $\lambda(M)$.
@@ 2.1. Extensions of valuations and Rees algebras
where $m_{v}$ is the maximal ideal of $\mathcal{O}_{v}$. Let $P=\mbox{center}_{R}(v)$. We have that $\sqrt{I(v)_{\lambda}\cap R}=P$ for all $\lambda\in vK_{>0}$ and in fact $I(v)_{\lambda}\cap R$ is $P$-primary, since the inclusion $R\rightarrow\mathcal{O}_{v}$ factors through $R_{P}$ and $m_{v}\cap R_{P}=PR_{P}$.
@@
Let $K\rightarrow L$ be a finite field extension, $R$ be a normal Noetherian domain with quotient field $K$ and $S$ be the integral closure of $R$ in $L$. Let $v$ be a valuation of $K$ which is nonnegative on $R$, and $w$ be an extension of $v$ to $L$. Then $w$ is nonnegative on $S$ and
@@
The maximal ideal $m_{w}$ of $\mathcal{O}_{w}$ contracts to the maximal ideal $m_{v}$ of $\mathcal{O}_{v}$ so $\mbox{center}_{S}(w)$ contracts to $\mbox{center}_{R}(v)$. Thus the lemma holds since $S$ is integral over $R$. ∎
For ideals $U$ and $V$ of a ring $R$, define
$U:V^{\infty}=\cup_{n>0}U:V^{n}=\{f\in R\mid fV^{n}\subset U\mbox{ for some $n>0$}\}.$
If $R$ is Noetherian, $U:V^{\infty}$ is obtained from $U$ by removing from a primary decomposition of $U$ the $P$-primary components that have the property that $J(R)\subset P$.
Let $R$ be a normal Noetherian semi local domain with quotient field $K$, $v_{1},\ldots,v_{r}$ be rank 1 valuations of $K$ which are nonnegative on $R$ and $\lambda_{i}\in v_{i}K_{\geq 0}$ for $1\leq i\leq r$. Let $J(R)$ be the intersection of the maximal ideals of $R$.
We have that
$\left(\cap_{i=1}^{r}I(v_{i})_{\lambda_{i}}\right)\cap R=\cap_{i=1}^{r}(I(v_{i})_{\lambda_{i}}\cap R).$
Let $P_{j}$ for $1\leq j\leq s$ be the distinct centers of the $v_{i}$ on $R$, and define $\sigma(i)$ by $\mbox{center}_{R}(v_{i})=P_{\sigma(i)}$ for $1\leq i\leq r$. As observed before Lemma 2.1, $I(v_{i})_{\lambda_{i}}\cap R$ is $P_{\sigma(i)}$-primary for all $i$. Let $Q_{j}=\cap_{\sigma(i)=j}(I(v_{i})_{\lambda_{i}}\cap R)$, which is a $P_{j}$-primary ideal since it is an intersection of $P_{j}$-primary ideals. Thus $\cap_{j=1}^{s}Q_{j}$ is a primary decomposition of $\cap_{i=1}^{r}(I(v_{i})_{\lambda_{i}}\cap R)$. As remarked above, $(\cap_{j=1}^{s}Q_{j}):J(R)^{\infty}$ removes from $\cap_{j=1}^{s}$ the primary components from the maximal ideals of $R$.
Thus
(10) $\left[\cap_{i\in\Lambda}I(v_{i})_{\lambda_{i}}\right]\cap R=\left(\left[\cap_{i=1}^{r}I(v_{i})_{\lambda_{i}}\right]\right)\cap R:J(R)^{\infty},$
@@
The following lemma follows from Lemma 2.1, Lemma 2.2 and equation (10).
@@
Let $K$ be a field and $v_{1},\ldots,v_{r}$ be rank 1 valuations of $K$ and $\lambda_{i}\geq 0$ be elements of $v_{i}K$ for $1\leq i\leq r$. Let $R$ be a normal semilocal domain with quotient field $K$ such that the $v_{i}$ are nonnegative on $R$ for all $i$. Define a graded filtration $\mathcal{I}(R)=\{I(R)_{n}\}$ of $R$ by $I(R)_{n}=(\cap_{i}I(v_{i})_{n\lambda_{i}})\cap R$. Suppose that $K\rightarrow L$ is a finite field extension. Let $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$ for $1\leq i\leq r$. Let $S$ be the integral closure of $R$ in $L$. Define a graded filtration
(11) $\mathcal{I}(S)=\{I(S)_{n}\}$ on $S$ by $I(S)_{n}=(\cap_{i}\cap_{j}I(w_{ij})_{n\lambda_{i}})\cap S$.
@@
Let $N$ be a finite extension of $L$ such that $N$ is a normal extension of $K$. Let $G=\mbox{Aut}_{K}(N)$. We have a tower of fields $K\rightarrow N^{G}\rightarrow N$ where $N^{G}$ is the fixed field of $N$ under the action of $G$. The extension $K\rightarrow N^{G}$ is purely inseparable and $N^{G}\rightarrow N$ is Galois with Galois group $G$ (c.f. [27, Proposition V.6.11]).
@@
(14) $w_{i,1}\sigma_{j}=w_{i,\tau_{i}(j)}$
for all $j$ ([42, Corollary VI.7.3]). For $b\in\{1,\ldots,a\}$, we compute
@@
where the minimum is over $\alpha_{1}<\alpha_{2}<\cdots<\alpha_{b}$. Thus
@@ 2.2. Divisorial Filtrations
A divisorial valuation of a Noetherian domain $R$ ([39, Definition 9.3.1]) is a valuation $\nu$ of $K$ such that if $\mathcal{O}_{\nu}$ is the valuation ring of $\nu$ with maximal ideal $\mathfrak{m}_{\nu}$, then $R\subset\mathcal{O}_{\nu}$ and if $\mathfrak{p}=\mathfrak{m}_{\nu}\cap R$ then $\mbox{trdeg}_{\varkappa(\mathfrak{p})}\varkappa(\nu)={\rm ht}(\mathfrak{p})-1$, where $\varkappa(\mathfrak{p})$ is the residue field of $R_{\mathfrak{p}}$ and $\varkappa(\nu)$ is the residue field of $\mathcal{O}_{\nu}$. If $R$ is a local ring and $\nu$ is a divisorial valuation of $R$ such that $m_{R}=m_{\nu}\cap R$, then $\nu$ is called an $m_{R}$-valuation.
@@
By [39, Theorem 9.3.2], the valuation ring of every divisorial valuation $\nu$ is Noetherian, hence is a discrete valuation. Suppose that $R$ is an excellent local domain. Then a valuation $\nu$ of the quotient field $K$ of $R$ which is nonnegative on $R$ is a divisorial valuation of $R$ if and only if the valuation ring $\mathcal{O}_{\nu}$ is essentially of finite type over $R$ ([13, Lemma 6.1] or [39, Theorem 9.3.2]).
Let $R$ be a Noetherian domain and $v$ be a discrete valuation of the quotient field of $R$ which is nonnegative on $R$. For $\lambda\in{\mathbb{R}}_{\geq 0}$, define the valuation ideal of $R$
$I_{R}(v)_{\lambda}=\{f\in R\mid v(f)\geq\lambda\}=I(v)_{\lambda}\cap R.$
A divisorial filtration of a Noetherian domain $R$ is a filtration =$\mathcal{I}=\{I_{n}\}$ where there exist divisorial valuations $v_{1},\ldots,v_{r}$ of $R$ and $\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{R}}_{\geq 0}$ such that
$I_{n}=I_{R}(v_{1})_{n\lambda_{1}}\cap\cdots\cap I_{R}(v_{r})_{n\lambda_{r}}=I(v_{1})_{n\lambda_{1}}\cap\cdots\cap I(v_{r})_{n\lambda_{r}}\cap R.$
We now weaken our assumptions on $R$ and suppose that $R$ is an arbitrary Noetherian local ring.
Definition 2.5.
Let $R$ be a Noetherian local ring and $P_{1},\ldots,P_{t}$ be the minimal prime ideals of $R$. Let $\varphi_{i}:R\rightarrow R/P_{i}$ be the natural projection for $1\leq i\leq t$. Let $\{v_{ij}\}$ be divisorial valuations on $R/P_{i}$ and $\lambda_{ij}\in{\mathbb{R}}_{\geq 0}$. Let =$\mathcal{I}=\{I_{n}\}$ where
(16) $I_{n}=\cap_{i=1}^{t}\cap_{j}\varphi_{i}^{-1}(I_{R/P_{i}}(v_{ij})_{n\lambda_{ij}}).$
If $P$ is a prime ideal of a local ring $R$, $\varphi:R\rightarrow R/P$ and $v$ is a valuation of the quotient field of $R/P$, then $w=v\circ\varphi$ is a valuation of $R$, as defined in VI.3.1 [2]. That is, by extending the value group $vR/P$ of $V$ by introducing the element $\infty$, which is larger than any element of $vR/P$, and letting $\Gamma_{\infty}=vR/P\cup\{\infty\}$, the mapping $w:R\rightarrow\Gamma_{\infty}$ satisfies the properties
(17) $\begin{array}[]{l}w(xy)=w(x)+w(y)\mbox{ for }x,y\in R\\ w(x+y)\geq\min\{w(x),w(y)\}\mbox{ for }x,y\in R\\ w(1)=0\mbox{ and }w(P)=\infty\end{array}$
With the notation of (16), the mappings $v_{ij}\circ\varphi_{i}$ are valuations of $R$, satisfying the conclusions of (17).
@@
Associated to a discrete valuation ring $V$ is a unique canonical valuation $v_{V}$ defined by $v_{V}(x)=n$ if $x\in m_{V}^{n}\setminus m_{V}^{n+1}$ for $x\in V$ and where $m_{V}$ is the maximal ideal of $V$. If $w$ is any valuation whose valuation ring is $V$, so that $w$ is an equivalent valuation to $v$, we have that $w=cv$ where $c=w(u)$ for $u$ a generator of $m_{V}$ (a uniformizing parameter of $V$).
Now suppose that $R$ is a normal excellent local domain. Let
(18) $\pi:X\rightarrow\mbox{Spec}(R)$
be a birational projective morphism such that $X$ is normal. Let
(19) $D=\sum\lambda_{i}E_{i}$
be an effective ${\mathbb{R}}$-Weil divisor on $X$ (so that all $\lambda_{i}\in{\mathbb{R}}_{\geq 0}$), where the $E_{i}$ are prime divisors on $X$. Define
$\lceil nD\rceil=\sum\lceil n\lambda_{i}\rceil E_{i},$
where $\lceil x\rceil$ is the round up of a real number $x$. Then for all $n$, $I(nD)=\Gamma(X,\mathcal{O}_{X}(-\lceil nD\rceil))$ is an ideal of $R$ and =$\mathcal{I}=\{I(nD)\}$ is a divisorial filtration of $R$.
Let $v_{1},\ldots,v_{r}$ be the canonical valuations of the valuation rings $\mathcal{O}_{X,E_{i}}$. Then
(20) $\Gamma(X,\mathcal{O}_{X}(-\lceil nD\rceil))=I_{R}(v_{1})_{n\lambda_{1}}\cap\cdots\cap I_{R}(v_{r})_{n\lambda_{r}}.$
Let =$\mathcal{I}=\{I_{n}\}$ be a divisorial filtration of $R$. By [13, Remark 6.6 to Lemma 5.5], there exists $\pi:X\rightarrow\mbox{Spec}(R)$ and $D$ as in (18) and (19) such that $I_{n}=I(nD)$ for all $n\geq 0$.
A choice of $X$ and $D$ such that $I_{n}=I(nD)$ for all $n\geq 0$ is called a representation of the divisorial filtration $\mathcal{I}$. The choice of $X$ is not unique, and the choice of $D$ on a particular $X$ may not be unique. It is possible for a particular $\mathcal{I}$ to have representations (20) which are ${\mathbb{Q}}$-divisorial and also have representations (20) which are not ${\mathbb{Q}}$-divisorial.
@@
Let $\{w_{ij}\}$ be the extensions of $v_{i}$ to $L$. Let $u_{i,j}$ be the canonical valuation of $\mathcal{O}_{w_{ij}}$, that is, a uniformizing parameter $t$ of $\mathcal{O}_{w_{i,j}}$ satisfies $u_{i,j}=1$. There exists a prime divisor $F_{ij}$ on $Y$ such that $\mathcal{O}_{w_{ij}}=\mathcal{O}_{F_{ij}}$. Write $w_{ij}=d_{ij}u_{ij}$. Let $D^{\prime}=\sum_{i,j}\frac{\lambda_{i}}{d_{ij}}F_{ij}$. Then
(21) $I(D^{\prime})_{n}=\cap_{i}\cap_{j}I(u_{ij})_{n\frac{\lambda_{i}}{d_{ij}}}\cap S=\cap_{i}\cap_{j}I(w_{ij})_{n\lambda_{i}}\cap S$
@@ 3. A generalized McAdam theorem
In this section we prove a theorem (Theorem 3.5) which is the essential part of the proof of Theorem 1.5 stated in the introduction.
@@ 3.1. Resolutions and Alterations
The proof of [12, Section 5] shows the following.
@@
([12, Section 5]) Let $R$ be a normal excellent local domain and let =$\mathcal{I}=\{I_{n}\}$ be a ${\mathbb{Q}}$-divisorial filtration on $R$. Suppose that there exists a projective birational morphism $\pi:X\rightarrow\mbox{Spec}(R)$ such that $X$ is nonsingular, and there exists an effective ${\mathbb{Q}}$-divisor $D$ on $X$ such that $\mathcal{I}=\mathcal{I}(D)$. If $m_{R}\in\mbox{Ass}(R/I_{m_{0}})$ for some $m_{0}\in{\mathbb{Z}}_{>0}$, then the analytic spread of $\mathcal{I}$ is $\ell(\mathcal{I})=\dim R$.
The following proposition is an extension of of [12, Section 6 ], with essentially the same proof.
@@
([12, Section 6] Let $R\rightarrow S$ be a finite extension of reduced rings where $R$ is a Noetherian local ring of dimension $d$ and $S$ is a semi local ring. Let $\mathcal{J}=\{J_{n}\}$ be a graded filtration of $S$ and let =$\mathcal{I}=\{I_{n}\}$ be the graded filtration of $R$ defined by $I_{n}=J_{n}\cap R$. For $q$ a prime ideal of $S$, let $\mathcal{J}_{q}$ be the graded filtration $\mathcal{J}_{q}=\{(J_{n})_{q}\}$ of $S_{q}$. Suppose that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$ for some $m_{0}>0$. Then there exists a maximal ideal $q$ of $S$ such that $qS_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$. If we further have that $R$ and $S$ are equidimensional of dimension $d$, $\ell(\mathcal{J}_{q})=d$ and there exists a nonzero divisor of $R$ in $M=\mbox{Ann}_{R}(S/R)$, then $\ell(\mathcal{I})=d$.
@@
Let $\{\mathfrak{m}_{i}$} be the maximal ideals of $S$. Since we assume that $m_{R}\in\mbox{Ass}_{R}(R/I_{m_{0}})$, there exists $y\in R/I_{n_{0}}$ such that $m_{R}=\mbox{ann}_{R}(y)$. We have that $y\in S/J_{n}$ is nonzero since $R/I_{m_{0}}\rightarrow S/J_{m_{0}}$ is an inclusion. Thus $\mbox{Ann}_{S}(y)\neq S$. Since maximal elements in the set of annihilators of elements of $S/J_{m_{0}}$ are prime ideals (by [30, Theorem 6.1]), there exists a prime ideal $q$ in $S$ which contains $\mbox{Ann}_{S}(y)$ and is the annihilator of an element $z$ of $S/J_{m_{0}}$. We have that $q$ contains $m_{R}S$ and
@@
so $q$ is a maximal ideal of $S$ such that $q\in\mbox{Ass}_{S}(S/J_{m_{0}})$, and thus $qS_{q}\in\mbox{Ass}_{S_{q}}(S_{q}/(J_{m_{0}})_{q})$.
@@
This follows from [13, Remark 6.6 to Lemma 6.5]. ∎
@@
([24, Definition 2.20]) Let $X$ be a Noetherian integral scheme. An Alteration $X^{\prime}$ of $X$ is an integral scheme $X^{\prime}$ with a morphism $\varphi:X^{\prime}\rightarrow X$ which is dominant, proper and such that for some open $U\subset X$, $\varphi^{-1}(U)\rightarrow U$ is finite.
A scheme $X$ is nonsingular if $\mathcal{O}_{X,q}$ is a regular local ring for all $q\in X$. An alteration $\varphi:X^{\prime}\rightarrow X$ is regular if $X^{\prime}$ is nonsingular.
3.2. Associated primes and maximal analytic spread
In this subsection we prove the following theorem. Theorem 3.5 is proven for excellent equicharacteristic zero local domains in Theorem 1.3 [12].